# Relative_Constructibility_via_Restricted_ASMs.bib

### v1

@@ 2.2 Modifying the Satisfaction Relation
An obvious cause of the triviality identified in the preceding paragraph, is the lack of restriction on the transition function $\tau_{\mathfrak{M}}$ of an ASM $\mathfrak{M}$. Ideally, $\tau_{\mathfrak{M}}(s)$ should be describable is a way which only depends locally on itself and $s$. One formalisation of a “describable local property” applicable to first-order structures is the notion of a first-order property. A first-order formula $\phi$ over the signature of a structure $\mathfrak{A}$ is a first-order property of $\mathfrak{A}$ iff $\mathfrak{A}\models\phi$, where $\models$ is Tarski’s satisfaction relation. However, in our use case, the formula $\phi$ might need to refer to two different structures, so some slight modifications to the satisfaction relation are in order.
@@
Given an signature $\sigma$, define $2\cdot\sigma$ to be the disjoint union of $\sigma$ and itself. In other words, for each symbol $\ulcorner X\urcorner\in\sigma$ and each $i<2$, we choose a fresh symbol $\ulcorner X_{i}\urcorner$ and cast it to be of the same type and arity as $\ulcorner X\urcorner$. We then set
@@
  • $X^{(\mathfrak{A}_{0},\mathfrak{A}_{1})}:=a$ if $\ulcorner X\urcorner=\ulcorner a\urcorner$ for some $a\in A$,
  • $X^{(\mathfrak{A}_{0},\mathfrak{A}_{1})}:=Y^{\mathfrak{A}_{0}}$ if $\ulcorner X\urcorner=\ulcorner Y_{0}\urcorner$ for some $\ulcorner Y\ulcorner\in\sigma$, and
  • $X^{(\mathfrak{A}_{0},\mathfrak{A}_{1})}:=Y^{\mathfrak{A}_{1}}$ if $\ulcorner X\urcorner=\ulcorner Y_{1}\urcorner$ for some $\ulcorner Y\ulcorner\in\sigma$.
@@
    $(\mathfrak{A}_{0},\mathfrak{A}_{1})\models_{2}\phi\text{ iff }t_{1}^{(\mathfrak{A}_{0},\mathfrak{A}_{1})}=t_{2}^{(\mathfrak{A}_{0},\mathfrak{A}_{1})}\text{.}$
@@
    $(\mathfrak{A}_{0},\mathfrak{A}_{1})\models_{2}\phi\text{ iff }R^{(\mathfrak{A}_{0},\mathfrak{A}_{1})}(t_{1}^{(\mathfrak{A}_{0},\mathfrak{A}_{1})},\dots,t_{n}^{(\mathfrak{A}_{0},\mathfrak{A}_{1})})\text{.}$
@@
    $(\mathfrak{A}_{0},\mathfrak{A}_{1})\models_{2}\phi\text{ iff }\neg((\mathfrak{A}_{0},\mathfrak{A}_{1})\models_{2}\psi)\text{.}$
@@
    $(\mathfrak{A}_{0},\mathfrak{A}_{1})\models_{2}\phi\text{ iff }((\mathfrak{A}_{0},\mathfrak{A}_{1})\models_{2}\psi_{1}\wedge(\mathfrak{A}_{0},\mathfrak{A}_{1})\models_{2}\psi_{2})\text{.}$
@@
    $(\mathfrak{A}_{0},\mathfrak{A}_{1})\models_{2}\phi\text{ iff }\exists a\in A((\mathfrak{A}_{0},\mathfrak{A}_{1})\models_{2}\psi[x\mapsto a])\text{,}$
@@
Before adding more requirements to the transition function of an ASM, let us enforce that every state in an ASM $\mathfrak{M}$ must have the same base set $b(\mathfrak{M})$. This can be done without loss of generality in the original definition of an ASM, because we can extend the base set of each state in $\mathfrak{M}$ to be the union of the base sets of states in $\mathfrak{M}$, while arbitrarily extending the interpretation of function and relation symbols in the signature by each state.
We stipulate that the transition function $\tau_{\mathfrak{M}}$ of an ASM $\mathfrak{M}$ must be finitely describable over $2\cdot\sigma(\mathfrak{M})$. That is, there must exist a first-order sentence $\phi^{\tau}_{\mathfrak{M}}$ over $2\cdot\sigma(\mathfrak{M})$ such that for all $s_{1},s_{2}\in S(\mathfrak{M})$,
@@
It turns out that we can intentionally choose the set of states in an ASM to trivialise the definition of its transition function. To overcome this, we make it so that the set of states depends only on the base set and signature of the structures involved. What this means is, if $\mathfrak{M}_{1}$ and $\mathfrak{M}_{2}$ are ASMs, $s_{1}\in S(\mathfrak{M}_{1})$, $s_{2}\in S(\mathfrak{M}_{2})$ and $s_{1}$, $s_{2}$ share the same base set and signature, then $S(\mathfrak{M}_{1})=S(\mathfrak{M}_{2})$. To wit, we include $b(\mathfrak{M})$ as a primary datum of an ASM, and derive $S(\mathfrak{M})$ from $b(\mathfrak{M})$ and $\sigma(\mathfrak{M})$ as follows:
$S(\mathfrak{M}):=\{s:s\text{ is a first-order structure with base set }b(\mathfrak{M})\text{ and signature }\sigma(\mathfrak{M})\}\text{.}$
Now, there may be a nagging worry that we over-corrected, that our modified definition of an ASM is too restrictive. Indeed, this shortcoming becomes obvious when we try to define an ASM corresponding to a Turing machine.
@@
Consider a Turing machine $M$ with an $\omega$-length tape and $n$ states, for some $n<\omega$. An ASM $\mathfrak{M}$ representing $M$ naturally has $b(\mathfrak{M})$ include $\omega$ and the set of $M$’s states. A suitable transition function $\tau_{\mathfrak{M}}$ ought to speak of $M$’s states, and to do that it is imperative that each of these states identify uniquely with a term over $\sigma(\mathfrak{M})$. For example, we can have a constant symbol $\ulcorner c_{i}\urcorner\in\sigma(\mathfrak{M})$ uniquely represent the $i$-th state of $M$.
@@ 2.3 Theories with Constraints in Interpretation
Theories with constraints in interpretation (henceforth, TCIs) were conceived in [myself] as a convenient means of looking at generic objects produced by set-theoretic forcing. As TCIs are considered a generalisation of first-order structures, it is only natural that we adapt them for the purpose of constraining the possible states of an ASM.
@@
(Lau, [myself]) A first-order theory with constraints in interpretation (first-order TCI) — henceforth, just theory with constraints in interpretation (TCI) — is a tuple $(T,\sigma,\dot{\mathcal{U}},\vartheta)$, where
@@
  • $\vartheta$ is a function (the interpretation constraint map) with domain $\sigma\cup\{\dot{\mathcal{U}}\}$,
@@
(Lau, [myself]) Let $(T,\sigma,\dot{\mathcal{U}},\vartheta)$ be a TCI. We say $\mathcal{M}:=(U;\mathcal{I})\models^{*}(T,\sigma,\dot{\mathcal{U}},\vartheta)$ — or $\mathcal{M}$ models $(T,\sigma,\dot{\mathcal{U}},\vartheta)$ — iff all of the following holds:
@@
    • $\circ$
      if $\dot{X}$ is an $n$-ary relation symbol and $\vartheta(\dot{X})=(y,0)$, then $\mathcal{I}(\dot{X})\subset y\cap U^{n}$,
    • $\circ$
      if $\dot{X}$ is an $n$-ary relation symbol and $\vartheta(\dot{X})=(y,1)$, then $\mathcal{I}(\dot{X})=y\cap U^{n}$,
    • $\circ$
      if $\dot{X}$ is an $n$-ary function symbol and $\vartheta(\dot{X})=(y,0)$, then
@@
    • $\circ$
      if $\dot{X}$ is an $n$-ary function symbol and $\vartheta(\dot{X})=(y,1)$, then
@@
To specify an ASM $\mathfrak{M}$, we should also specify a TCI $\mathfrak{T}(\mathfrak{M})=(T,\sigma,\dot{\mathcal{U}},\vartheta)$ such that
@@
Let us now check that any Turing machine can be easily represented as an ASM.
@@
$I(\mathfrak{M}):=\{s:$ $s\models^{*}\mathfrak{T}(\mathfrak{M})\text{ and }$
@@
  • the read/write head at the 00-th cell of the tape,
@@
  • $1$’ written on exactly one cell of the tape, and
  • 00’ written on all other cells of the tape.
@@
We are almost done with specifying our ASM $\mathfrak{M}$; the only thing left to do is to give a transition function $\tau_{\mathfrak{M}}$ accurately representing transitions among the Turing configurations of $M$. It suffices to concoct a suitable sentence $\phi^{\tau}_{\mathfrak{M}}$ based on the transition function
@@
$\phi_{j,k}:=$ $\ulcorner(q(R_{0}(h_{0}),k)\wedge t_{0}\!=\!\langle j\rangle)\!\!\implies\!\!(q(R_{1}(h_{0}),k^{\prime})\wedge t_{1}\!=\!\langle j^{\prime}\rangle\wedge h_{1}\!=\mathrm{Pre}(h_{0})\wedge\varphi)\urcorner$
[…]
$\phi_{j,k}:=$ $\ulcorner(q(R_{0}(h_{0}),k)\wedge t_{0}\!=\!\langle j\rangle)\!\!\implies\!\!(q(R_{1}(h_{0}),k^{\prime})\wedge t_{1}\!=\!\langle j^{\prime}\rangle\wedge h_{1}\!=\mathrm{Suc}(h_{0})\wedge\varphi)\urcorner$
[…]
$\phi_{j,k}:=$ $\ulcorner h_{0}=h_{1}\wedge t_{0}=t_{1}\wedge\forall x\ (R_{0}(x)\iff R_{1}(x))\urcorner$
[…]
@@
2.4 Absoluteness and the Bounded Exploration Postulate
@@
Recall that it is implicit in our modified definition of an ASM $\mathfrak{M}$, that $\phi^{\tau}_{\mathfrak{M}}$ needs to have the following property:
@@
However, given any $\mathfrak{T}(\mathfrak{M})$ and $\phi^{\tau}_{\mathfrak{M}}$, it is not immediate that (2.1) is absolute for transitive models of $\mathsf{ZFC}$. As a result, it is difficult to ascertain that the definition of $\tau_{\mathfrak{M}}$ from $\mathfrak{T}(\mathfrak{M})$ and $\phi^{\tau}_{\mathfrak{M}}$ is absolute for transitive models of $\mathsf{ZFC}$. This implores us to make the dependency between the current and the next state of an ASM even more local.
At this point, it is useful to make precise Gurevich’s bounded exploration postulate, for it is an example of highly local dependency. Combined with our stipulation that a transition function must be definable, we expect to obtain a level of local definability from which the absoluteness we want can be easily derived. Consequently, abiding by the bounded exploration postulate should net us absoluteness at no additional cost.
@@
Let $\mathfrak{M}$ be an ASM and $s\in S(\mathfrak{M})$. Define $\Delta_{\mathfrak{M}}(s)$ to be the unique function from $\sigma(\mathfrak{M})$ into $\mathcal{P}(b(\mathfrak{M}))$ such that
@@
    $\Delta_{\mathfrak{M}}(s)(\ulcorner X\urcorner)=\{\vec{z}\in b(\mathfrak{M})^{n}:\ulcorner X\urcorner^{s}(\vec{z})\neq\ulcorner X\urcorner^{\tau_{\mathfrak{M}}(s)}(\vec{z})\}\text{,}$
@@
Definition 2.17.
Let $\mathfrak{M}$ be an ASM. Its transition function $\tau_{\mathfrak{M}}$ satisfies the bounded exploration postulate iff there is a finite set of ground terms $\mathcal{D}$ over $\sigma(\mathfrak{M})$ such that for all $\ulcorner X\urcorner\in\sigma(\mathfrak{M})$ and $s_{0},s_{1}\in S(\mathfrak{M})$,
@@
Consider ASMs $\mathfrak{M}_{1}$ and $\mathfrak{M}_{2}$ such that
@@
On the other hand, $\tau_{\mathfrak{M}_{2}}$ does not satisfy the bounded exploration postulate because for $s_{0},s_{1}\in S(\mathfrak{M}_{1})$, $\Delta_{\mathfrak{M}}(s_{0})=\Delta_{\mathfrak{M}}(s_{1})$ only if $\ulcorner\mathrm{In}\urcorner^{s_{0}}=\ulcorner\mathrm{In}\urcorner^{s_{1}}$, and whether $\ulcorner\mathrm{In}\urcorner^{s_{0}}=\ulcorner\mathrm{In}\urcorner^{s_{1}}$ cannot be determined by just examining a finite fragment of $s_{0}$ and a finite fragment of $s_{1}$. This means the transition function of any ASM built on $\tau_{\mathfrak{M}_{2}}$ by enlarging its signature while maintaining its function (“erase all data from the input tape and do nothing else”), cannot possibly satisfy the postulate.
@@
Besides the erasure of a tape’s contents, one would expect an ASM $\mathfrak{M}_{\text{copy}}$ which copies the input tape’s contents to the output tape, and does nothing else, to only carry out steps as valid as universal bit-flipping. Yet, by an argument similar to that for $\tau_{\mathfrak{M}_{2}}$ not satisfying the bounded exploration postulate in Example 2.18, $\mathfrak{M}_{\text{copy}}$ fails to satisfy the same postulate.
@@
[…]
@@
    Points (C2) and (C3) guarantee that whenever $\mathfrak{M}$ is an ASM with $\sigma(\mathfrak{M})=\sigma$, the statement
@@
    Suppose $\phi=\phi^{\tau}_{\mathfrak{M}}$ for some ASM $\mathfrak{M}$. Then in the context of $\mathfrak{M}$, for each $\ulcorner X\urcorner\in\sigma$, the value of the next state’s interpretation of $\ulcorner X\urcorner$ at a point depends only on finitely many values of the current state’s interpretations of symbols in $\sigma$. This agrees in spirit with the requirement of bounded exploration.

### v2

@@ 2.2 Modifying the Satisfaction Relation
An obvious cause of the triviality identified in Example 2.1, is the lack of restriction on the transition function $\tau_{\mathfrak{M}}$ of a SeqA $\mathfrak{M}$. Ideally, $\tau_{\mathfrak{M}}(s)$ should be describable is a way which only depends locally on itself and $s$. One formalisation of a “describable local property” applicable to first-order structures is the notion of a first-order property. A first-order formula $\phi$ over the signature of a structure $\mathfrak{A}$ is a first-order property of $\mathfrak{A}$ iff $\mathfrak{A}\models\phi$, where $\models$ is Tarski’s satisfaction relation. To describe a transition function, we want to isolate a first-order property of the next state that depends solely on the current state. However, because a candidate property should refer to two different structures (the current and next states), some slight modifications to the satisfaction relation are in order.
@@
Given a signature $\sigma$, define $2\cdot\sigma$ to be the disjoint union of $\sigma$ and itself. In other words, for each symbol $\ulcorner X\urcorner\in\sigma$ and each $i<2$, we choose a fresh symbol $\ulcorner X_{i}\urcorner$ and cast it to be of the same type and arity as $\ulcorner X\urcorner$. We then set
@@
$X^{(\mathfrak{A}_{0},\mathfrak{A}_{1})}:=\$$a\text{ if }\ulcorner X\urcorner=\ulcorner a\urcorner\text{ for some }a\in A\text{,}$
$X^{(\mathfrak{A}_{0},\mathfrak{A}_{1})}:=\$$Y^{\mathfrak{A}_{0}}\text{ if }\ulcorner X\urcorner=\ulcorner Y_{0}\urcorner\text{ for some }\ulcorner Y\urcorner\in\sigma\text{, and}$
$X^{(\mathfrak{A}_{0},\mathfrak{A}_{1})}:=\$$Y^{\mathfrak{A}_{1}}\text{ if }\ulcorner X\urcorner=\ulcorner Y_{1}\urcorner\text{ for some }\ulcorner Y\urcorner\in\sigma\text{.}$
@@
[…]
  2. (2)
@@
[…]
  3. (3)
@@
[…]
  4. (4)
@@
[…]
  5. (5)
@@
[…]
@@
Before adding more requirements to the transition function of a SeqA, let us enforce that every state in a SeqA $\mathfrak{M}$ must have the same base set $b(\mathfrak{M})$. This can be done without loss of generality in the definition of a SeqA, because we can extend the base set of each state in $\mathfrak{M}$ to be the union of the base sets of states in $\mathfrak{M}$, while arbitrarily extending the interpretation of function and relation symbols in the signature by each state.
We stipulate that the transition function $\tau_{\mathfrak{M}}$ of a SeqA $\mathfrak{M}$ must be finitely describable over $2\cdot\sigma(\mathfrak{M})$. That is, there must exist a first-order sentence $\phi^{\tau}_{\mathfrak{M}}$ over $2\cdot\sigma(\mathfrak{M})$ such that for all $s_{1},s_{2}\in S(\mathfrak{M})$,
@@
It turns out that we can intentionally choose the set of states in a SeqA to trivialise the definition of its transition function. To overcome this, we make it so that the set of states depends only on the base set and signature of the structures involved. What this means is, if $\mathfrak{M}_{1}$ and $\mathfrak{M}_{2}$ are SeqAs, $s_{1}\in S(\mathfrak{M}_{1})$, $s_{2}\in S(\mathfrak{M}_{2})$ and $s_{1}$, $s_{2}$ share the same base set and signature, then $S(\mathfrak{M}_{1})=S(\mathfrak{M}_{2})$. To wit, we include $b(\mathfrak{M})$ as a primary datum of a SeqA, and derive $S(\mathfrak{M})$ from $b(\mathfrak{M})$ and $\sigma(\mathfrak{M})$ as follows:
[…]
Now, there may be a nagging worry that we over-corrected, that our modified definition of a SeqA is too restrictive. Indeed, this shortcoming becomes obvious when we try to define a SeqA corresponding to a Turing machine.
@@
Consider a Turing machine $M$ with an $\omega$-length tape and $n$ states, for some $n<\omega$. A SeqA $\mathfrak{M}$ representing $M$ naturally has $b(\mathfrak{M})$ include $\omega$ and the set of $M$’s states. A suitable transition function $\tau_{\mathfrak{M}}$ ought to speak of $M$’s states, and to do that it is imperative that each of these states identify uniquely with a term over $\sigma(\mathfrak{M})$. For example, we can have a constant symbol $\ulcorner c_{i}\urcorner\in\sigma(\mathfrak{M})$ uniquely represent the $i$-th state of $M$.
@@ 2.3 Theories with Constraints in Interpretation
Theories with constraints in interpretation (henceforth, TCIs) were conceived as a convenient means of looking at generic objects produced by set-theoretic forcing. As TCIs impose natural constraints on first-order theories, they can be easily utilised to constrain the possible states of a SeqA. In fact, TCIs are a generalisation of what Dershowitz and Gurevich call static functions in [15].
@@
(Lau, [19]) A first-order theory with constraints in interpretation (first-order TCI) — henceforth, just theory with constraints in interpretation (TCI) — is a tuple $(T,\sigma,\dot{\mathcal{U}},\vartheta)$, where
  1. (a)
@@
  3. (c)
    $\vartheta$ is a function (the constraint function) with domain $\sigma\cup\{\dot{\mathcal{U}}\}$,
  4. (d)
@@
(Lau, [19]) Let $(T,\sigma,\dot{\mathcal{U}},\vartheta)$ be a TCI. We say
$\mathcal{M}:=(U;\mathcal{I})\models^{*}(T,\sigma,\dot{\mathcal{U}},\vartheta)$
— or $\mathcal{M}$ models $(T,\sigma,\dot{\mathcal{U}},\vartheta)$ — iff all of the following holds:
  1. (a)
@@
    2. (ii)
      if $\dot{X}$ is a $n$-ary relation symbol and $\vartheta(\dot{X})=(y,0)$, then $\mathcal{I}(\dot{X})\subset y\cap U^{n}$,
    3. (iii)
      if $\dot{X}$ is a $n$-ary relation symbol and $\vartheta(\dot{X})=(y,1)$, then $\mathcal{I}(\dot{X})=y\cap U^{n}$,
    4. (iv)
      if $\dot{X}$ is a $n$-ary function symbol and $\vartheta(\dot{X})=(y,0)$, then
@@
    5. (v)
      if $\dot{X}$ is a $n$-ary function symbol and $\vartheta(\dot{X})=(y,1)$, then
@@
To specify a SeqA $\mathfrak{M}$, we should also specify a TCI $\mathfrak{T}(\mathfrak{M})=(T,\sigma,\dot{\mathcal{U}},\vartheta)$ such that
@@
Let us now check that any Turing machine can be easily represented as a SeqA.
@@
$I(\mathfrak{M}):=\{s:\$$s\models^{*}\mathfrak{T}(\mathfrak{M})\text{ and }$
@@
  1. (a)
    the read/write head at the 0-th cell of the tape,
  2. (b)
@@
  3. (c)
    11’ written on exactly one cell of the tape, and
  4. (d)
    0’ written on all other cells of the tape.
@@
We are almost done with specifying our SeqA $\mathfrak{M}$; the only thing left to do is to give a transition function $\tau_{\mathfrak{M}}$ accurately representing transitions among the Turing configurations of $M$. It suffices to concoct a suitable sentence $\phi^{\tau}_{\mathfrak{M}}$ based on the transition function
@@
$\phi_{j,k}:=\$$\ulcorner(q(R_{0}(h_{0}),k)\wedge t_{0}\!=\!\langle j\rangle)\!\!\implies\!\!(q(R_{1}(h_{0}),k^{\prime})\wedge t_{1}\!=\!\langle j^{\prime}\rangle\wedge h_{1}\!=\mathrm{Pre}(h_{0})\wedge\varphi)\urcorner$
$\text{ if }j<n-1\text{ and }\delta((c_{j},k))=(c_{j^{\prime}},k^{\prime},\mathrm{left})\text{,}$
$\phi_{j,k}:=\$$\ulcorner(q(R_{0}(h_{0}),k)\wedge t_{0}\!=\!\langle j\rangle)\!\!\implies\!\!(q(R_{1}(h_{0}),k^{\prime})\wedge t_{1}\!=\!\langle j^{\prime}\rangle\wedge h_{1}\!=\mathrm{Suc}(h_{0})\wedge\varphi)\urcorner$
$\text{ if }j<n-1\text{ and }\delta((c_{j},k))=(c_{j^{\prime}},k^{\prime},\mathrm{right})\text{, and}$
$\phi_{j,k}:=\$$\ulcorner h_{0}=h_{1}\wedge t_{0}=t_{1}\wedge\forall x\ (R_{0}(x)\iff R_{1}(x))\urcorner$
$\text{ if }j=n-1\text{,}$
@@
2.4 Absoluteness and the Bounded Exploration Postulate Revisited
@@
Recall that it is implicit in our modified definition of a SeqA $\mathfrak{M}$, that $\phi^{\tau}_{\mathfrak{M}}$ needs to have the following property:
@@
However, given any $\mathfrak{T}(\mathfrak{M})$ and $\phi^{\tau}_{\mathfrak{M}}$, it is not immediate that (2.1) is absolute for transitive models of $\mathsf{ZFC}$. As a result, it is difficult to ascertain that the definition of $\tau_{\mathfrak{M}}$ from $\mathfrak{T}(\mathfrak{M})$ and $\phi^{\tau}_{\mathfrak{M}}$ is absolute for transitive models of $\mathsf{ZFC}$. This implores us to make the dependency between the current and the next state of a SeqA even more local.
At this point, it is useful to revisit, in a more precise fashion, Gurevich’s bounded exploration postulate, for it is an example of highly local dependency even without the closure of the state space under isomorphisms. Combined with our stipulation that a transition function must be definable, we expect to obtain a level of local definability from which the absoluteness we want can be easily derived. Consequently, abiding by the bounded exploration postulate should net us absoluteness at no additional cost.
@@
Let $\mathfrak{M}$ be a SeqA and $s\in S(\mathfrak{M})$. Define $\Delta_{\mathfrak{M}}(s)$ to be the unique function from $\sigma(\mathfrak{M})$ into $\mathcal{P}(b(\mathfrak{M}))$ such that
  1. (a)
@@
    $\Delta_{\mathfrak{M}}(s)(\ulcorner X\urcorner)=\{\vec{z}^{\frown}(c)\in b(\mathfrak{M})^{n+1}:\ulcorner X\urcorner^{s}(\vec{z})\neq\ulcorner X\urcorner^{\tau_{\mathfrak{M}}(s)}(\vec{z})=c\}\text{,}$
  2. (b)
@@
Definition 2.17 (Gurevich, [7]).
Let $\mathfrak{M}$ be a SeqA. Its transition function $\tau_{\mathfrak{M}}$ satisfies the bounded exploration postulate iff there is a finite set of ground terms $\mathcal{D}$ over $\sigma(\mathfrak{M})$ such that for all $\ulcorner X\urcorner\in\sigma(\mathfrak{M})$ and $s_{0},s_{1}\in S(\mathfrak{M})$,
@@
Consider SeqAs $\mathfrak{M}_{1}$ and $\mathfrak{M}_{2}$ such that
@@
On the other hand, $\tau_{\mathfrak{M}_{2}}$ does not satisfy the bounded exploration postulate because for $s_{0},s_{1}\in S(\mathfrak{M}_{1})$, $\Delta_{\mathfrak{M}}(s_{0})=\Delta_{\mathfrak{M}}(s_{1})$ only if $\ulcorner\mathrm{In}\urcorner^{s_{0}}=\ulcorner\mathrm{In}\urcorner^{s_{1}}$, and whether $\ulcorner\mathrm{In}\urcorner^{s_{0}}=\ulcorner\mathrm{In}\urcorner^{s_{1}}$ cannot be determined by just examining a finite fragment of $s_{0}$ and a finite fragment of $s_{1}$. This means the transition function of any SeqA built on $\tau_{\mathfrak{M}_{2}}$ by enlarging its signature while maintaining its function (“erase all data from the input tape and do nothing else”), cannot possibly satisfy the postulate.
@@
Besides the erasure of a tape’s contents, one would expect a SeqA $\mathfrak{M}_{\text{copy}}$ which copies the input tape’s contents to the output tape, and does nothing else, to only carry out steps as valid as universal bit-flipping. Yet, by an argument similar to that for $\tau_{\mathfrak{M}_{2}}$ not satisfying the bounded exploration postulate in Example 2.18, $\mathfrak{M}_{\text{copy}}$ fails to satisfy the same postulate.
@@
    If there is no structure $\mathfrak{B}$ with signature $\sigma$ and base set $A$, for which $(\mathfrak{A},\mathfrak{B})\models_{2}\phi$, then the cause of said non-existence is necessarily witnessed by a $\Delta_{0}$-definable (i.e. definable by a first-order formula over $\{\in\}$ with only bounded quantifiers) fragment of $\mathfrak{A}$’s partial valuation map of one of the $\phi^{\ulcorner X\urcorner}$s, with parameters in $A^{<\omega}$. As $\Delta_{0}$ definitions are absolute for transitive models of $\mathsf{ZFC}$, the existence of a witness to the non-existence such $\mathfrak{B}$ must be too.
@@
    Points (C2) and (C3) guarantee that whenever $\mathfrak{M}$ is a SeqA with $\sigma(\mathfrak{M})=\sigma$, the statement
@@
    Suppose $\phi=\phi^{\tau}_{\mathfrak{M}}$ for some SeqA $\mathfrak{M}$. Then in the context of $\mathfrak{M}$, for each $\ulcorner X\urcorner\in\sigma$, the value of the next state’s interpretation of $\ulcorner X\urcorner$ at a point depends only on finitely many values of the current state’s interpretations of symbols in $\sigma$. This agrees in spirit with the requirement of bounded exploration, albeit only locally. The caveat of bounded exploration being local is a reasonable — even expected — compromise for generalised computation.